OpenStudy (anonymous):

rrange these numbers from least to greatest: 5/6, 7/9, 3/5 A. 5/6, 3/5, 7/9 B. 5/6, 7/9, 3/5 C. 3/5, 7/9, 5/6 D. 3/5, 5/6, 7/9

OpenStudy (anonymous):

it is a

OpenStudy (anonymous):

Mb its C

OpenStudy (anonymous):

3/5 = 0.6 7/9= 0.78 5/6= 0.83

OpenStudy (anonymous):

actually im a ma'am

OpenStudy (anonymous):

GG.

OpenStudy (anonymous):

Lol sorry ma'am.

OpenStudy (anonymous):

XDD thanks guys

OpenStudy (anonymous):

Np Friend, even though you copied mah epic face. D:

OpenStudy (anonymous):

nooo, mines green, and says words lol XP ill change it l8r